OpenStudy (anonymous):

Solve.. (3x-1)^2-2(3x-1) I know the answer is 9x^2-12x+3 But, I don't see why..

OpenStudy (anonymous):

substitue (3x-1) with say x adn solve the quadratic

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i mean simplyfy... so it would become x^2-2x... which becomes x(x-2) ... then substitute back in (3x-1) and simplyfy

OpenStudy (campbell_st):

ok... just expand or distribute each part \[(3x - 1)^2 = 9x^2 - 6x + 1\] and \[2(3x - 1) = 6x - 2\] finding the difference gives \[(3x -1)^2 - 2(3x -1) = 9x^2 - 6x + 1 -(6x - 2)\] all you need to do is collect the like terms

OpenStudy (campbell_st):

@aceace I think the question should have been simplify.... as this is an expression so it can't be solved.

OpenStudy (anonymous):

yeh i rectified that in one of my replies

OpenStudy (anonymous):

just factor it